A father is three times as old as his son, and his daughter is 3 years younger than the son.If the sum if their ages 3 years ago was 63 years, find the present age of the father.
Answer by addingup(3677)

f = 3s
d = s-3
f-3+s-3+d-3 = 63 substitute for f and d per above equalities:
3s-3+s-3+s-3-3 = 63
5s-12 =63
5s = 75
s = 15 This is how old the son is.
f = 3(15)= 45
d = 15-3 = 12